12. Select the correct type of system represented by the following system:
6x + 2y = 10
3x + y = 5
A. Consistent and independent
B. Inconsistent
C. Consistent and dependent
D. None of these


Answer by ewatrrr(24785)


Hi
6x + 2y = 10
3x + y = 5
6x + 2y = 10
-6x- 2y = -10
0 = 0 this is a dependent system with infinitely many solutions
Represent the 'same' Line when graphed
A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution.
A dependent system has infinitely many solutions.